# Current Trends in Digital Twin Development, Maintenance, and Operation: An Interview Study

### v3

@@ 6.1.4 Discussion
Based on our findings presented in this section, there is no common understanding of DTs across the nineteen interviewees. It is relevant to mention that this lack of common understanding has been discussed as one of the non-technical challenges in DT development by van den Brand et al. [28] and it is highly important to overcome this. Moreover, we aimed to understand the alignment of interviewees on the different parts that make up a DT, especially the highly relevant ones such as the relation and communication between AE and virtual counterpart. It can also be observed from our findings that interviewees did not agree on all the components of a DT.
@@ 6.2 Influence of reuse on DT lifecycle (RQ2)
As indicated by Walravens et al. [42], developing DTs is a cross-domain and resource intensive task where reuse of existing artifacts can significantly reduce the development and maintenance costs of DTs. With RQ2 our objective was to gain insight into the existing practices of reusing various artifacts for DT development. We identified 15 out of 19 interviewees acknowledging some form of artifact reuse while working with DTs. Industrial participants mentioned reuse more frequently. Based on the interviews, we identified two kinds of artifact reuse: data reuse and software component reuse, which we discuss in Section 6.2.1 and summarize in Table 5. We also discuss several related challenges that restrict the possibilities of reuse in Section 6.2.2. Afterwards, in Section 6.2.3 we discuss our interpretation of these findings and what insights can be derived from them.
@@ 6.2.1 Current practices of reusing artifacts
  • Reduced resources for development: Four interviewees mentioned that reusing software artifacts lead to shorter delivery time and reduced development effort. According to them, specially the ones from industry, reuse is essential as it greatly affects time to market. We additionally identified two cases where DTs were developed based on one or more existing DTs, allowing the developers to leverage existing artifacts and reduce development effort significantly. These identified benefits confirm our assumptions on the benefits of software artifact reuse based on prior publications reporting similar benefits for more traditional software systems [17, 6].
  • Ease of use: According to six interviewees, ease of use and built-in support for integration provided by corresponding tooling encouraged them to reuse software artifacts. Based on our analysis, we divided these tools and integrated environments into two categories: commercial and in-house tools. While the commercially available tools are used both in industry and academia, the tools built in-house are exclusively used by the corresponding companies. Among the commercially available tools, Unity [50], a well-known physics-based game engine, stood out as six interviewees mentioned to have used it or one of its commercial extension for developing geometry and physics models. Furthermore, our analysis suggests that in-house tools are often developed based on requirements defined by the organizations themselves and therefore are only suitable for their specific needs.
@@ 6.3 Consistency management of cross-domain models (RQ3)
Cross-domain collaboration is essential for the development and maintenance of DTs. Within a cross-domain environment, we identified maintaining consistency among cross-domain software models as a challenge [28]. With RQ3 we aimed to investigate this challenge. We wanted to understand how consistency is defined in practice; and identify key causes for inconsistencies, and tools and techniques used to manage them. Twelve interviewees mentioned that they have encountered or put measures in place to handle inconsistencies. In the following sections we present our findings based on the analysis of the data we collected from these interviewees.
@@ 6.3.1 Inconsistencies encountered in practice
  • Interface inconsistency: Hisarciklilar et al. [10] defined interface inconsistency as mismatching values, terminologies, or schemes among connected interface elements. We identified this inconsistency mostly in cases where two or more models, often cross-domain, need to communicate and do not share any compatible interfaces. Our analysis suggests that this is one of the most frequently encountered inconsistencies, with four interviewees explicitly mentioning it.
@@
  • Design-implementation inconsistency: As DTs often represent complex cyber-physical systems, various diagrams (i.e., UML [48], SysML [46]) are used to conceptually represent parts or the complete system, usually during the design phase. In our analysis we identified cases where the actual implementation deviates from the design, which we identify as representation inconsistency.
@@ 6.3.4 Discussion
Our analysis suggests that inconsistencies are actual issues in the context of DTs and directly or indirectly affected over 60% of our interviewees. The nature and source of these issues are highly diverse and depend on the actual DT implementation, involved methods and tooling, organizational and personal constraints. Consequently, we believe that it is fairly impossible to categorise these inconsistencies completely. Furthermore, we identified that inconsistency issues are often not categorized as such. Instead, they are treated as regular issues encountered during system development or maintenance. Therefore, we believe there is a lack of awareness of the need for specialized methodologies or techniques for identifying and mitigating inconsistencies. We think this might have contributed to the inadequacy of related tooling, as discussed before. Furthermore, we identified situations where the vast majority of inconsistency issues are being avoided by simply using a single tool for development, or by developing large monolithic models. Besides, we think there is a large gap between academic innovation and industrial practice in the context of inconsistency management. For example, Torres et al. [25] referred to several academically developed consistency management approaches in their systematic literature review. However, we could not identify any commonality between this list and the approaches discussed by our interviewees, which we present in Section 6.3.3. Therefore, we believe that there are opportunities for further research and development for consistency management tools in the context of DTs. However, we also identified that certain domains, i.e., aerospace, construction engineering, automotive, are more aware of and mature about handling inconsistency issues. This is largely facilitated by safety requirements and standards or conventions accepted across related organizational entities.
@@ 6.4.1 Model integration practices
  • Multi-tool approach. This approach is used when a heterogeneous modeling environment is present, in which distinct modeling tools are combined [49]. Each model needs encapsulation resulting in a defined interface to establish communication with other models. The nine interviewees using the multi-tool approach agree that this approach has advantages for cross-domain collaboration because it allows different tools to be used. Other reasons mentioned to use this approach is information hiding of model details also known as model masking, for IP protection and reduction of model complexity. Here, technologies are used so the model becomes a “black box” with only its input and output exposed, and hiding the model details. However, this approach has many challenges such as interface consistency (discussed in Section 6.3.1), data and model semantics, and relationship complexity between the models.
@@
  • FMI: Three interviewees specifically mentioned the use of FMI technology to integrate their models. This technology has been used in DTs [20] and is supported by over 170 modeling tools [^4] https://fmi-standard.org/. According to three interviewees from industry, two reasons to use this technology are its maturity and compatibility with various modeling tools.
@@ 6.5 Model orchestration (RQ5)
We define DT model orchestration as the definition of the required communication actions and correct execution sequence of the models [5]. To achieve this the orchestration should include activities such as data compatibility checks and starting and ending model execution [9]. With RQ5 we aimed to understand interviewees’ perceptions and practices related to such orchestration in DTs.
@@ 6.5.1 Model orchestration practices
  • Commercial. We identified three types of commercial tools used as execution platforms. The first is to use modeling tools to orchestrate the models, e.g., MATLAB. This type of tool is used by two interviewees who also use the single-tool approach for the integration of models as explained in Section 6.4.1. The second type is the use of system design tools such as IBM Rhapsody [47] and HEEDS [^5] https://www.plm.automation.siemens.com/global/en/products/simcenter/simcenter-heeds.html, used by two interviewees. These tools facilitate the use of external software as long as they are supported by the vendor. The third type encompasses tools that support a DSL to sequence entities for execution; this tool type is used by two interviewees. The tools mentioned by interviewees are PDDL (Planning Domain Definition Language) and Dezyne from Verum [^6] https://www.verum.com/DiscoverDezyne. The latter also supports formal verification.
@@ 6.6.1 Verification and validation techniques
    Moreover, we identified that deeper observational tests are done by creating visual representations such as graphs or 3D visuals of the behavior of both AE and VE and superimposing them to observe the extent of overlapping and differences. In other cases, AE and VE behaviors are translated into events and actions in a Gantt chart [1]. The timing and sequence of actions are compared between AE and VE to check if there are any differences. When there is a combination of continuous and discrete behaviors, continuous signals are transformed into discrete ones and then the behaviors of AE and VE are compared to check for equivalence.
@@
  • Formal methods and tools: We identified two interviewees from industry and one interviewee from academia who mentioned using formal methods in DT development. Formal verification techniques have been used to validate the behavior of DTs with the help of tools such as Verum’s Dezyne [27] and Coco [^7] https://cocotec.io/. An interviewee from industry mentioned that DSLs have been used to specify the behavior of a system and transform such specifications into timed automata models in UPPAAL [2], allowing model checking. However, model checking is not always scalable enough, considering state space explosion [7]. The interviewee discussed that in order to address this challenge, recurring behavioral patterns were identified and validated using model checking, rather than the entire system. In this way, some level of correctness guarantees was provided. Another interviewee mentioned that formal methods have been used for consistent execution of models in DTs: formal semantics for such execution were defined and were helpful to understand differences in execution between models. He also shared that for their model interfaces, they formally proved that the components adhere to the interfaces to avoid interface violations caused by component changes. He mentioned that this ensured consistent behavior when integrating components. As witnessed by the above, three interviewees mentioned formal techniques for validating DTs; no others did. We speculate that the lower adoption of this method could be attributed to scalability issues.
    As mentioned before, formal techniques such as model checking are not scalable, due to state space explosion problems [7]. In addition, we believe that the complexity of formal methods and engineers’ lack of background in them, could possibly contribute to the lower adoption of these techniques.
@@
Apart from these two categories of techniques identified and quantitatively analysed by us, DTs also often undergo testing across their entire lifecycle, in order to check adherence to requirements. Interviewees used the term ‘testing’ to discuss two different items, namely, using scripts to test the system and comparing the AE and VE behavior using observational tests. Due to this lack of clarity, we did not perform quantitative analysis of interviewee responses for testing. As mentioned earlier, in some domains such as the space domain, testing is highly time and resource intensive. In such cases, careful consideration is needed on when and what aspects to test, based on the DT and its context. Moreover, in these cases, testing effort is then balanced with effectiveness. For instance, after resolving an integration issue, only local tests are done. On the other hand, full regression testing would be executed when replacing an entire sub-component. An observation worth mentioning here is that this is not only specific to testing DTs, but generally used in the context of testing software systems. Different types of tests have been mentioned by interviewees such as model-based testing, integration testing, and unit testing. In some cases, static analysis is also performed to detect coding errors, thus, helping gain confidence about the system. Some tools mentioned by interviewees are Axini’s [^8] https://www.axini.com/en/, Matlab [16], and Unity [50] which are used for creating and testing models on the fly.
@@ 6.6.2 Challenges in validation of DTs
Apart from the challenges discussed in the interviews, we identified that the continuous evolution of VEs pose an additional challenge for validation of VEs. This is because of its continuous synchronization with AE, and others as discussed by Zhang et al. [31], which requires specific techniques for continuous validation of VE as it keeps evolving [28].
@@ 6.7.1 Properties for validation of DTs
Fidelity of VE
8 5
Time difference in execution between AE and VE
2 3
Related to DT’s purpose, application, or domain
9 6
Related to dynamic consistency
5 2
@@ 6.7.3 Discussion
From our findings, we noticed that some interviewees expressed their concerns on how to measure the quality of a DT and how to quantify the relevant properties. We also observed that works such as Dalibor et al. [34] discuss this concern on quality assurance and requirements for DTs. Our interviewees, from a range of domains, uniformly agreed that the properties for DT validation depend on the DT’s domain, purpose, and application.
@@ 7.6 Discussion on additional findings
Takeaway message: Our analysis shows that DT architectures should facilitate maintainability, cross-domain collaboration, and rapid development. It also shows that DT development processes contain ingredients of software development processes or adaptations thereof for DTs. A common recommendation for development is to define the goal of the DT before starting its development. Predictions on the business model of DTs have been made such as DTs and their related data becoming a tradeable asset whose ownership could be transferred or shared across the lifecycle. DTs have been predicted to improve the automation and self-adaptability of systems; and also to help in the design support for such systems. Current challenges in DT development such as lack of intuitiveness and standards are predicted to be resolved in the future.
@@ 8 Threats to validity
This research is an empirical study, which is never completely devoid of omissions or pitfalls. We discuss the threats to validity (TTV) as described by Easterbrook et al. [8] and the measures taken to mitigate these.
@@ 9 Related work
This paper conducts a study to identify categories of barriers hindering the development of DTs in the process industry. The paper proposes a model designed to recognize potential facilitators addressing the identified barriers and intends to offer possible solutions.
SLR
Segovia et al. [39]
@@
This paper analyses the state-of-practice on DT development, with a specific focus on the methodological approaches found in scientific literature. The paper considers three primary phases: design, modeling, and implementation. It states the challenges that are encountered in each phase.
SLR
Sharma et al. [40]
@@
This study focuses on the conceptualization and adoption of DT in practical applications. The goal of this research is to define the key components of DTs and identify open research questions for their development. This study analyses three case studies; two from the construction domain and one from the medical domain.
SLR
Liu et al. [35]
@@
This paper performs an analysis of the approaches to develop DT based on a classification. Specifically, its implementation classification criteria are based on a central component. For instance, if the primary component is a service, the DT is categorized as a ”DT as a service.
SLR
Trauer et al. [41]
@@
The objective of this paper is to investigate the obstacles faced by 61 practitioners in implementing DTs. It particularly aims to understand the characteristics of DTs and the technical and non-technical challenges associated with their implementation.
Interview
Mihai et al. [36]
@@
This research aims to understand the technologies that ease the implementation of DT and its challenges. It identifies enabling technologies and outlines research challenges for DT implementation, including fidelity and synchronization of models and ongoing standardization efforts.
SLR
Botin-Sanabria et al. [33]
@@
This paper examines different barriers to the use and development of DTs, revealing high-level obstacles spanning social, technical, and business domains.
SLR
Parnianifard et al. [37]
@@
This research discusses limitations and future challenges, highlighting the complexity and interoperability issues arising from the heterogeneous nature of DT components.
SLR
Ammar et al. [32]
@@
This study aims to understand the usage and challenges of DTs by interviewing eight practitioners in the construction industry. Six categories of challenges are identified, with a focus on technology development, particularly in sensor technology.
Interview
@@
We have identified two main research topics from the papers presented in Table 13. The first domain is associated with methodologies employed in the implementation DTs [39, 40, 35, 32]. Research related with methodologies for implementing DTs has yielded valuable outputs by identifying potential open research questions and approaches embraced by practitioners. However, it is noteworthy that the focus of these investigations predominantly revolves around the primary approaches by practitioners. This differs from our research where we focus our effort on specific development areas, such as cross-domain model consistency, orchestration, and dynamic consistency.
@@
The second topic revolves around the identification of challenges and enablers in the implementation of DTs [38, 41, 36, 33, 37]. Each of these works has highlighted diverse technical and non-technical challenges confronted by practitioners in the course of DT implementation. Nevertheless, these studies avoid examining deeply the strategies employed by practitioners in addressing these challenges. Our research focused particularly on exploring the strategies used by practitioners to deal with different implementation challenges.

### v4

@@ 6.1.4 Discussion
Based on our findings presented in this section, there is no common understanding of DTs across the nineteen interviewees. It is relevant to mention that this lack of common understanding has been discussed as one of the non-technical challenges in DT development by van den Brand et al. [Models_meet_data] and it is highly important to overcome this. Moreover, we aimed to understand the alignment of interviewees on the different parts that make up a DT, especially the highly relevant ones such as the relation and communication between AE and virtual counterpart. It can also be observed from our findings that interviewees did not agree on all the components of a DT.
@@ 6.2 Influence of reuse on DT lifecycle (RQ2)
As indicated by Walravens et al. [walravensTurtle], developing DTs is a cross-domain and resource intensive task where reuse of existing artifacts can significantly reduce the development and maintenance costs of DTs. With RQ2 our objective was to gain insight into the existing practices of reusing various artifacts for DT development. We identified 15 out of 19 interviewees acknowledging some form of artifact reuse while working with DTs. Industrial participants mentioned reuse more frequently. Based on the interviews, we identified two kinds of artifact reuse: data reuse and software component reuse, which we discuss in Section 6.2.1 and summarize in Table 5. We also discuss several related challenges that restrict the possibilities of reuse in Section 6.2.2. Afterwards, in Section 6.2.3 we discuss our interpretation of these findings and what insights can be derived from them.
@@ 6.2.1 Current practices of reusing artifacts
  • Reduced resources for development: Four interviewees mentioned that reusing software artifacts lead to shorter delivery time and reduced development effort. According to them, specially the ones from industry, reuse is essential as it greatly affects time to market. We additionally identified two cases where DTs were developed based on one or more existing DTs, allowing the developers to leverage existing artifacts and reduce development effort significantly. These identified benefits confirm our assumptions on the benefits of software artifact reuse based on prior publications reporting similar benefits for more traditional software systems [BARROSJUSTO20181, Mohagheghi2007].
  • Ease of use: According to six interviewees, ease of use and built-in support for integration provided by corresponding tooling encouraged them to reuse software artifacts. Based on our analysis, we divided these tools and integrated environments into two categories: commercial and in-house tools. While the commercially available tools are used both in industry and academia, the tools built in-house are exclusively used by the corresponding companies. Among the commercially available tools, Unity [Unity_ref], a well-known physics-based game engine, stood out as six interviewees mentioned to have used it or one of its commercial extension for developing geometry and physics models. Furthermore, our analysis suggests that in-house tools are often developed based on requirements defined by the organizations themselves and therefore are only suitable for their specific needs.
@@ 6.3 Consistency management of cross-domain models (RQ3)
Cross-domain collaboration is essential for the development and maintenance of DTs. Within a cross-domain environment, we identified maintaining consistency among cross-domain software models as a challenge [Models_meet_data]. With RQ3 we aimed to investigate this challenge. We wanted to understand how consistency is defined in practice; and identify key causes for inconsistencies, and tools and techniques used to manage them. Twelve interviewees mentioned that they have encountered or put measures in place to handle inconsistencies. In the following sections we present our findings based on the analysis of the data we collected from these interviewees.
@@ 6.3.1 Inconsistencies encountered in practice
  • Interface inconsistency: Hisarciklilar et al. [Hisarciklilar2011] defined interface inconsistency as mismatching values, terminologies, or schemes among connected interface elements. We identified this inconsistency mostly in cases where two or more models, often cross-domain, need to communicate and do not share any compatible interfaces. Our analysis suggests that this is one of the most frequently encountered inconsistencies, with four interviewees explicitly mentioning it.
@@
  • Design-implementation inconsistency: As DTs often represent complex cyber-physical systems, various diagrams (i.e., UML [om2017unified], SysML [hause2006sysml]) are used to conceptually represent parts or the complete system, usually during the design phase. In our analysis we identified cases where the actual implementation deviates from the design, which we identify as representation inconsistency.
@@ 6.3.4 Discussion
Our analysis suggests that inconsistencies are actual issues in the context of DTs and directly or indirectly affected over 60% of our interviewees. The nature and source of these issues are highly diverse and depend on the actual DT implementation, involved methods and tooling, organizational and personal constraints. Consequently, we believe that it is fairly impossible to categorise these inconsistencies completely. Furthermore, we identified that inconsistency issues are often not categorized as such. Instead, they are treated as regular issues encountered during system development or maintenance. Therefore, we believe there is a lack of awareness of the need for specialized methodologies or techniques for identifying and mitigating inconsistencies. We think this might have contributed to the inadequacy of related tooling, as discussed before. Furthermore, we identified situations where the vast majority of inconsistency issues are being avoided by simply using a single tool for development, or by developing large monolithic models. Besides, we think there is a large gap between academic innovation and industrial practice in the context of inconsistency management. For example, Torres et al. [torres2020systematic] referred to several academically developed consistency management approaches in their systematic literature review. However, we could not identify any commonality between this list and the approaches discussed by our interviewees, which we present in Section 6.3.3. Therefore, we believe that there are opportunities for further research and development for consistency management tools in the context of DTs. However, we also identified that certain domains, i.e., aerospace, construction engineering, automotive, are more aware of and mature about handling inconsistency issues. This is largely facilitated by safety requirements and standards or conventions accepted across related organizational entities.
@@ 6.4.1 Model integration practices
  • Multi-tool approach. This approach is used when a heterogeneous modeling environment is present, in which distinct modeling tools are combined [Ptolemaeus:14:SystemDesign]. Each model needs encapsulation resulting in a defined interface to establish communication with other models. The nine interviewees using the multi-tool approach agree that this approach has advantages for cross-domain collaboration because it allows different tools to be used. Other reasons mentioned to use this approach is information hiding of model details also known as model masking, for IP protection and reduction of model complexity. Here, technologies are used so the model becomes a “black box” with only its input and output exposed, and hiding the model details. However, this approach has many challenges such as interface consistency (discussed in Section 6.3.1), data and model semantics, and relationship complexity between the models.
@@
  • FMI: Three interviewees specifically mentioned the use of FMI technology to integrate their models. This technology has been used in DTs [FMU_in_DTs] and is supported by over 170 modeling tools [^4] https://fmi-standard.org/. According to three interviewees from industry, two reasons to use this technology are its maturity and compatibility with various modeling tools.
@@ 6.5 Model orchestration (RQ5)
We define DT model orchestration as the definition of the required communication actions and correct execution sequence of the models [Standard_orchestration]. To achieve this the orchestration should include activities such as data compatibility checks and starting and ending model execution [BP_heterogeneousSys]. With RQ5 we aimed to understand interviewees’ perceptions and practices related to such orchestration in DTs.
@@ 6.5.1 Model orchestration practices
  • Commercial. We identified three types of commercial tools used as execution platforms. The first is to use modeling tools to orchestrate the models, e.g., MATLAB. This type of tool is used by two interviewees who also use the single-tool approach for the integration of models as explained in Section 6.4.1. The second type is the use of system design tools such as IBM Rhapsody [IBMUserGuide] and HEEDS [^5] https://www.plm.automation.siemens.com/global/en/products/simcenter/simcenter-heeds.html, used by two interviewees. These tools facilitate the use of external software as long as they are supported by the vendor. The third type encompasses tools that support a DSL to sequence entities for execution; this tool type is used by two interviewees. The tools mentioned by interviewees are PDDL (Planning Domain Definition Language) and Dezyne from Verum [^6] https://www.verum.com/DiscoverDezyne. The latter also supports formal verification.
@@ 6.6.1 Verification and validation techniques
    Moreover, we identified that deeper observational tests are done by creating visual representations such as graphs or 3D visuals of the behavior of both AE and VE and superimposing them to observe the extent of overlapping and differences. In other cases, AE and VE behaviors are translated into events and actions in a Gantt chart [wilson2003gantt]. The timing and sequence of actions are compared between AE and VE to check if there are any differences. When there is a combination of continuous and discrete behaviors, continuous signals are transformed into discrete ones and then the behaviors of AE and VE are compared to check for equivalence.
@@
  • Formal methods and tools: We identified two interviewees from industry and one interviewee from academia who mentioned using formal methods in DT development. Formal verification techniques have been used to validate the behavior of DTs with the help of tools such as Verum’s Dezyne [van2021dezyne] and Coco [^7] https://cocotec.io/. An interviewee from industry mentioned that DSLs have been used to specify the behavior of a system and transform such specifications into timed automata models in UPPAAL [behrmann2004tutorial], allowing model checking. However, model checking is not always scalable enough, considering state space explosion [baier2008principles]. The interviewee discussed that in order to address this challenge, recurring behavioral patterns were identified and validated using model checking, rather than the entire system. In this way, some level of correctness guarantees was provided. Another interviewee mentioned that formal methods have been used for consistent execution of models in DTs: formal semantics for such execution were defined and were helpful to understand differences in execution between models. He also shared that for their model interfaces, they formally proved that the components adhere to the interfaces to avoid interface violations caused by component changes. He mentioned that this ensured consistent behavior when integrating components. As witnessed by the above, three interviewees mentioned formal techniques for validating DTs; no others did. We speculate that the lower adoption of this method could be attributed to scalability issues.
    As mentioned before, formal techniques such as model checking are not scalable, due to state space explosion problems [baier2008principles]. In addition, we believe that the complexity of formal methods and engineers’ lack of background in them, could possibly contribute to the lower adoption of these techniques.
@@
Apart from these two categories of techniques identified and quantitatively analysed by us, DTs also often undergo testing across their entire lifecycle, in order to check adherence to requirements. Interviewees used the term ‘testing’ to discuss two different items, namely, using scripts to test the system and comparing the AE and VE behavior using observational tests. Due to this lack of clarity, we did not perform quantitative analysis of interviewee responses for testing. As mentioned earlier, in some domains such as the space domain, testing is highly time and resource intensive. In such cases, careful consideration is needed on when and what aspects to test, based on the DT and its context. Moreover, in these cases, testing effort is then balanced with effectiveness. For instance, after resolving an integration issue, only local tests are done. On the other hand, full regression testing would be executed when replacing an entire sub-component. An observation worth mentioning here is that this is not only specific to testing DTs, but generally used in the context of testing software systems. Different types of tests have been mentioned by interviewees such as model-based testing, integration testing, and unit testing. In some cases, static analysis is also performed to detect coding errors, thus, helping gain confidence about the system. Some tools mentioned by interviewees are Axini’s [^8] https://www.axini.com/en/, Matlab [higham2016matlab], and Unity [Unity_ref] which are used for creating and testing models on the fly.
@@ 6.6.2 Challenges in validation of DTs
Apart from the challenges discussed in the interviews, we identified that the continuous evolution of VEs pose an additional challenge for validation of VEs. This is because of its continuous synchronization with AE, and others as discussed by Zhang et al. [ZHANG2021151], which requires specific techniques for continuous validation of VE as it keeps evolving [Models_meet_data].
@@ 6.7.1 Properties for validation of DTs
Fidelity of VE 8 5
Time difference in execution between AE and VE 2 3
Related to DT’s purpose, application, or domain 9 6
Related to dynamic consistency 5 2
@@ 6.7.3 Discussion
From our findings, we noticed that some interviewees expressed their concerns on how to measure the quality of a DT and how to quantify the relevant properties. We also observed that works such as Dalibor et al. [Cross-Domain_SystMapping] discuss this concern on quality assurance and requirements for DTs. Our interviewees, from a range of domains, uniformly agreed that the properties for DT validation depend on the DT’s domain, purpose, and application.
@@ 7.6 Discussion on additional findings
Takeaway message:
Our analysis shows that DT architectures should facilitate maintainability, cross-domain collaboration, and rapid development. It also shows that DT development processes contain ingredients of software development processes or adaptations thereof for DTs. A common recommendation for development is to define the goal of the DT before starting its development. Predictions on the business model of DTs have been made such as DTs and their related data becoming a tradeable asset whose ownership could be transferred or shared across the lifecycle. DTs have been predicted to improve the automation and self-adaptability of systems; and also to help in the design support for such systems. Current challenges in DT development such as lack of intuitiveness and standards are predicted to be resolved in the future.
@@ 8 Threats to validity
This research is an empirical study, which is never completely devoid of omissions or pitfalls. We discuss the threats to validity (TTV) as described by Easterbrook et al. [easterbrook2008selecting] and the measures taken to mitigate these.
@@ 9 Related work
Perno et al. [RW_ProcessInds] This paper conducts a study to identify categories of barriers hindering the development of DTs in the process industry. The paper proposes a model designed to recognize potential facilitators addressing the identified barriers and intends to offer possible solutions. SLR
@@
Segovia et al. [RW_Model_imp] This paper analyses the state-of-practice on DT development, with a specific focus on the methodological approaches found in scientific literature. The paper considers three primary phases: design, modeling, and implementation. It states the challenges that are encountered in each phase. SLR
@@
Sharma et al. [DT_state_art_1] This study focuses on the conceptualization and adoption of DT in practical applications. The goal of this research is to define the key components of DTs and identify open research questions for their development. This study analyses three case studies; two from the construction domain and one from the medical domain. SLR
@@
Liu et al. [RW_SoA_survey] This paper performs an analysis of the approaches to develop DT based on a classification. Specifically, its implementation classification criteria are based on a central component. For instance, if the primary component is a service, the DT is categorized as a ”DT as a service. SLR
@@
Trauer et al. [RW_ImpChallenge] The objective of this paper is to investigate the obstacles faced by 61 practitioners in implementing DTs. It particularly aims to understand the characteristics of DTs and the technical and non-technical challenges associated with their implementation. Interview
@@
Mihai et al. [9899718] This research aims to understand the technologies that ease the implementation of DT and its challenges. It identifies enabling technologies and outlines research challenges for DT implementation, including fidelity and synchronization of models and ongoing standardization efforts. SLR
@@
Botin-Sanabria et al. [RW_challengeApp] This paper examines different barriers to the use and development of DTs, revealing high-level obstacles spanning social, technical, and business domains. SLR
@@
Parnianifard et al. [RW_survey_DTandCPS] This research discusses limitations and future challenges, highlighting the complexity and interoperability issues arising from the heterogeneous nature of DT components. SLR
@@
Ammar et al. [RW_DT_construction] This study aims to understand the usage and challenges of DTs by interviewing eight practitioners in the construction industry. Six categories of challenges are identified, with a focus on technology development, particularly in sensor technology. Interview
@@
We have identified two main research topics from the papers presented in Table 13. The first domain is associated with methodologies employed in the implementation DTs [RW_Model_imp, DT_state_art_1, RW_SoA_survey, RW_DT_construction]. Research related with methodologies for implementing DTs has yielded valuable outputs by identifying potential open research questions and approaches embraced by practitioners. However, it is noteworthy that the focus of these investigations predominantly revolves around the primary approaches by practitioners. This differs from our research where we focus our effort on specific development areas, such as cross-domain model consistency, orchestration, and dynamic consistency.
@@
The second topic revolves around the identification of challenges and enablers in the implementation of DTs [RW_ProcessInds, RW_ImpChallenge, 9899718, RW_challengeApp, RW_survey_DTandCPS]. Each of these works has highlighted diverse technical and non-technical challenges confronted by practitioners in the course of DT implementation. Nevertheless, these studies avoid examining deeply the strategies employed by practitioners in addressing these challenges. Our research focused particularly on exploring the strategies used by practitioners to deal with different implementation challenges.